if the area of triangle TIN is 26cm squared and its base is 6 cm, find the height of the triangle.
Answer by checkley79(3341)

AREA=BH/2
26=6H/2 CROSS MULTIPLY.
6H=26*2
6H=52
H=52/6
H=8.667 ANS. FOR THE HEIGHT.
PROOF:
26=6*8.667/2
26=52/2
26=26